OpenStudy (anonymous):

Find the sixth term of an arithmetic sequence with t1 = 2 and tn = tn-1 + 4.

OpenStudy (anonymous):

t6=6-1+4 t6=9

OpenStudy (anonymous):

that's not the answer though....

OpenStudy (anonymous):

\[t_{6}=t _{5}+4\] since we need to add 4 to get the next term 2,6,10,14,18 \[t _{6}=18\] by formular\[t _{6}=14+4=18\]

OpenStudy (lgbasallote):

@Jonask i'd hate to be the bearer of bad news....but 18 is the 5th term NOT the 6th term...you need to add 4 more to get the 6th term

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

thanks lgbasallote for the clarity

OpenStudy (lgbasallote):

just doing my job as the avatar

OpenStudy (anonymous):

doing a good job